# OntoChat: a Framework for Conversational Ontology Engineering using Language Models

### v1

@@ 1 Introduction
Ontology Engineering (OE) encompasses a number of activities defining a collaborative effort to design, evaluate, and reuse ontologies of general or domain-specific purposes [19, 25]. Depending on the intended reuse by one or more stakeholders, the process starts with ontology designers eliciting requirements from the former, while interacting with experts to validate their formal understanding of the domain. Although the way OE activities are organised may vary depending on the methodology adopted [9, 26], these interactions are frequent and iterative throughout the development cycle. In large projects, ontology designers engage in a number of manual activities and interact with multiple parties to make sure requirements are collected consistently and comprehensively, seeking clarification from domain experts to formalise and test them [11]. This multi-party interaction may require substantial resources and create systematic ambiguities and biases arising from potentially conflicting or ill-formulated requirements. Not only does this directly affect the ontology design and evaluation, but it may also jeopardise their reuse by the same stakeholders. In other cases, a blurry line exists among stakeholders, domain experts, and ontology designers, which is common to multidisciplinary and highly collaborative projects, such as Wikidata [29].
@@
  • Evidence on which OE activities are most in need of computational support, gathered from a survey involving participants with OE expertise.
@@ 3.5 Implementation details
OntoChat is implemented in Python 3.11 and is released on GitHub  [^2] https://github.com/King-s-Knowledge-Graph-Lab/OntoChat under the MIT license. To facilitate its use and collect user feedback, we implemented an interface prototype using Gradio [^3] https://www.gradio.app. This can be launched on a local server, also hosted on Hugging Face Spaces [^4] https://huggingface.co/spaces/b289zhan/OntoChat. The interface has four tabs that wrap all the aforementioned functionalities within the same environment. The current version uses OpenAI’s API and has been evaluated on the GPT-3 family of models (gpt-3.5-turbo, and gpt-3.5-turbo-16k for larger contexts).
@@ 4.1 Experimental methodology
Given the OWL definition of Music Meta, and the full set of (manually annotated) CQ driving its implementation, we evaluate this component as a classification task. By extending this set with adversarial CQs, we expect OntoChat to correctly discriminate between these groups.
@@ 4.2 Preliminary results
Finally, by extending the original 28 CQs with the same number of negative CQs (expressing requirements that are not yet supported by Music Meta), we found that OntoChat can correctly classify supported requirements with an accuracy of 87.5% (P=88%, R=85.7%). More details are provided in Appx. 0.B.4.
@@ 5 Conclusions
Ontology engineering in large collaborative projects faces challenges in managing complex interactions between stakeholders, domain experts, and ontology designers, leading to potential bottlenecks and inconsistencies in the resulting ontologies. After collecting strong evidence on the activities that are the most in need of computational support, we designed OntoChat, a framework implementing a conversational approach to ontology engineering, leveraging LLMs for requirement elicitation, analysis, and ontology testing. Evaluation results indicate positive reception among both domain experts and ontology engineers, with the framework seen as a tool that could significantly speed up traditional ontology engineering activities. Nonetheless, areas for improvement were also identified, such as enhancing the generation of data examples in user stories, refining name entity scope when creating CQs, and extending support for CQ analysis. While future work will focus on improving the framework based on the evaluation feedback and conducting a more comprehensive evaluation, this research underscores the potential of LLMs and conversational interfaces to enhance efficiency and collaboration in ontology engineering.
@@ 0.C.4 Competency question clustering
CQ
Which award was received by a music artist?
Test prediction
Yes
Expected test outcome
Yes (supported)
Explanation
The ontology includes the class ”Award” and the relation ”Received award” which connects a MusicArtist to the Award they have received. This allows for querying the ontology to find the award(s) received by a specific music artist.
CQ
In which time interval did the creation process took place?
@@
Expected test outcome
Yes (supported)
Explanation
@@
The ontology described can address the competency question ”In which time interval did the creation process take place?” The ontology includes classes and properties related to the creative process, such as CreativeProcess, CreativeAction, and AnnotationMethod. It also includes properties like ”isDerivedFrom” and ”Creates” that can be used to track the creation process and infer the time interval in which it took place.
CQ
Which is the recording process that recorded a musical performance?
Test prediction
Yes
Expected test outcome
Yes (supported)
Explanation
@@
The ontology can address the competency question ”Which is the recording process that recorded a musical performance?” The ontology includes the class ”RecordingProcess” and the property ”isRecordedBy” which connects a RecordingProcess to a MusicEntity or MusicalPerformance.
CQ
Does a music algorithm favor a specific genre?
Test prediction
Yes
Expected test outcome
No (not supported, but potentially plausible)
Explanation
@@
The ontology includes the necessary classes and relations to represent music algorithms, music genres, and the relationship between them. By using the ”isTrainedOn” and ”isTrainedBy” relations, it is possible to infer whether a music algorithm has been trained on a specific music genre or if a music genre has influenced the training of a music algorithm.
CQ
Is a music work associated to any case of plagiarism?
Test prediction
No
Expected test outcome
No (not supported)
Explanation
@@
The ontology focuses on describing music metadata related to artists, compositions, performances, and recordings, but it does not include specific concepts or relations related to plagiarism detection or identifying cases of plagiarism in music works. To address the competency question, additional concepts and relations related to plagiarism detection and analysis would need to be added to the ontology.
CQ
Which language is most used in a music artist’s lyrics?
Test prediction
No
Expected test outcome
No (not supported)
Explanation
@@
The ontology described does not include any classes or properties related to language or lyrics analysis. It focuses more on the structural and organisational aspects of music metadata, such as compositions, performances, recordings, and their relationships. To address the question of which language is most used in a music artist’s lyrics, additional classes and properties related to language analysis and text processing would need to be included in the ontology.
CQ
When was the album first sold?
Test prediction
Yes
Expected test outcome
No (not supported)
Explanation
@@
The ontology includes classes and properties related to albums, such as Release, Publication Situation, and Publisher. By using these classes and properties, it is possible to model the publication and release of albums, including the specific situation or context in which they are made available to the public. The ontology can capture information about the timing, location, and associated entities involved in the publication process, allowing for the inference of when the album was first sold.

### v2

@@ 1 Introduction
Ontology Engineering (OE) encompasses a number of activities defining a collaborative effort to design, evaluate, and reuse ontologies of general or domain-specific purposes [20, 26]. Depending on the intended reuse by one or more stakeholders, the process starts with ontology designers eliciting requirements from the former, while interacting with experts to validate their formal understanding of the domain. Although the way OE activities are organised may vary depending on the methodology adopted [10, 27], these interactions are frequent and iterative throughout the development cycle. In large projects, ontology designers engage in a number of manual activities and interact with multiple parties to make sure requirements are collected consistently and comprehensively, seeking clarification from domain experts to formalise and test them [12]. This multi-party interaction may require substantial resources and create systematic ambiguities and biases arising from potentially conflicting or ill-formulated requirements [6]. Not only does this affect the ontology design and evaluation, but it may also jeopardise their reuse by the same stakeholders. In other cases, a blurry line exists among stakeholders, domain experts, and ontology designers, which is common to multidisciplinary and highly collaborative projects, such as Wikidata [30].
@@
  • An investigation of which OE activities are most in need of computational support, gathered from a survey involving participants with OE expertise.
@@ 3.5 Implementation details
OntoChat is implemented in Python 3.11 and is released on GitHub [^2] https://github.com/King-s-Knowledge-Graph-Lab/OntoChat (code, prompts, experiments) under the MIT license. To facilitate its use and collect user feedback, we implemented an interface prototype using Gradio [1]. This can be launched on a local server, also hosted on Hugging Face Spaces [^3] https://huggingface.co/spaces/b289zhan/OntoChat. The interface has four tabs that wrap all the functionalities within the same environment. The current version uses OpenAI’s API and has been evaluated on the GPT-3 family of models (gpt-3.5-turbo, and gpt-3.5-turbo-16k for larger contexts).
@@ 4.1 Experimental methodology
Given the OWL definition of Music Meta, and the 28 (manually produced) CQs driving its implementation, we evaluate this component as a classification task. This is done by extending the CQ set with the same number of negative CQs (requirements that are not yet supported by Music Meta). We expect OntoChat to correctly discriminate between these groups.
@@ 4.2 Preliminary results
Finally, for ontology testing, we found that OntoChat can correctly classify supported requirements with an accuracy of 87.5% (P=88%, R=85.7%), and often provides explanations and examples to support its prediction (Appx. 0.B.4).
@@ 5 Conclusions
This work addresses the challenges of ontology engineering in large collaborative projects by implementing a conversational workflow to streamline the process. The proposed framework, OntoChat, leverages LLMs to facilitate requirement elicitation, analysis, and ontology testing. Our preliminary evaluation efforts demonstrate a positive response from domain experts and ontology engineers, indicating potential for accelerating conventional ontology engineering tasks.
Nonetheless, several limitations still exist, notably those inherent to the use of LLMs in specialised domains due to their limited or potentially obsolete knowledge. Additional challenges include addressing biases in persona creation and enhancing the framework to provide insights into implementation costs and timelines. This will allow us to measure the amount of user supervision and involvement (e.g., number of interactions with the LLM, specificity of user feedback) during the refinement steps, needed to achieve a reasonable output from OntoChat (e.g., a user story, a list of competency questions), in contrast to full manual curation. Future work will focus on addressing these challenges, while enhancing the generation of examples in user stories, refining named entity scope in competency question creation, and broadening analysis support.
5.0.1 Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101004746. This work was partly funded by the HE project MuseIT, which has been co-founded by the European Union under the Grant Agreement No 101061441.
@@ 0.C.4 Ontology testing
CQ Which award was received by a music artist?
Test prediction Yes
Expected test outcome Yes (supported)
Explanation The ontology includes the class ”Award” and the relation ”Received award” which connects a MusicArtist to the Award they have received. This allows for querying the ontology to find the award(s) received by a specific music artist.
CQ In which time interval did the creation process took place?
@@
Expected test outcome Yes (supported)
@@
Explanation The ontology described can address the competency question ”In which time interval did the creation process take place?” The ontology includes classes and properties related to the creative process, such as CreativeProcess, CreativeAction, and AnnotationMethod. It also includes properties like ”isDerivedFrom” and ”Creates” that can be used to track the creation process and infer the time interval in which it took place.
CQ Which is the recording process that recorded a musical performance?
Test prediction Yes
Expected test outcome Yes (supported)
@@
Explanation The ontology can address the competency question ”Which is the recording process that recorded a musical performance?” The ontology includes the class ”RecordingProcess” and the property ”isRecordedBy” which connects a RecordingProcess to a MusicEntity or MusicalPerformance.
CQ Does a music algorithm favor a specific genre?
Test prediction Yes
Expected test outcome No (not supported, but potentially plausible)
@@
Explanation The ontology includes the necessary classes and relations to represent music algorithms, music genres, and the relationship between them. By using the ”isTrainedOn” and ”isTrainedBy” relations, it is possible to infer whether a music algorithm has been trained on a specific music genre or if a music genre has influenced the training of a music algorithm.
CQ Is a music work associated to any case of plagiarism?
Test prediction No
Expected test outcome No (not supported)
@@
Explanation The ontology focuses on describing music metadata related to artists, compositions, performances, and recordings, but it does not include specific concepts or relations related to plagiarism detection or identifying cases of plagiarism in music works. To address the competency question, additional concepts and relations related to plagiarism detection and analysis would need to be added to the ontology.
CQ Which language is most used in a music artist’s lyrics?
Test prediction No
Expected test outcome No (not supported)
@@
Explanation The ontology described does not include any classes or properties related to language or lyrics analysis. It focuses more on the structural and organisational aspects of music metadata, such as compositions, performances, recordings, and their relationships. To address the question of which language is most used in a music artist’s lyrics, additional classes and properties related to language analysis and text processing would need to be included in the ontology.
CQ When was the album first sold?
Test prediction Yes
Expected test outcome No (not supported)
@@
Explanation The ontology includes classes and properties related to albums, such as Release, Publication Situation, and Publisher. By using these classes and properties, it is possible to model the publication and release of albums, including the specific situation or context in which they are made available to the public. The ontology can capture information about the timing, location, and associated entities involved in the publication process, allowing for the inference of when the album was first sold.